OpenStudy (anonymous):

use the elmination method to solve the system of equations a+7b=15 a+4b=6 There is a solution, it is ______ There is no solution There are infinetly many solutions

OpenStudy (tyteen4a03):

Subtract the first equation from the second equation to get 3b = 9, then continue solving.

OpenStudy (anonymous):

so then I would plug what I got into the first equation, 7(3) +a=15 which woud be 21+15 equals 6, so my ordered pair would be (3,6) right?

OpenStudy (tyteen4a03):

Correct. Note that your order was (b, a), not sure if you want (a, b) or not

OpenStudy (anonymous):

oh, thank you for pointing that out so it would be (6,3)

OpenStudy (tyteen4a03):

Oh wait, 7(3) +a=15 would give you a = 15-21, which would not be 6 (but -6 instead)

OpenStudy (anonymous):

ooh, now I am starting to really understand the elimination method.